OpenStudy (anonymous):

Evaluate: 8 + g ÷ h • 4, when g = 9 and h = 3.

razor99 (razor99):

ohh srry it should be like this 8+9/3x4

OpenStudy (anonymous):

17/12

razor99 (razor99):

=8+3x4

razor99 (razor99):

=8+12 =20

razor99 (razor99):

thats should be the answer

OpenStudy (unklerhaukus):

I am not sure if i have interpreted you question correctly where are the brackets \[8+g \div h \times 4\]\[8+{4g\over h}\]\[g=9,h=3\]\[\rightarrow 8+{4\times 9\over 3}\]\[=8+{36\over 12}\]\[=8+3\]\[=11\]

OpenStudy (unklerhaukus):

\[8 + (g ÷ h) \times 4≠8 + g ÷ (h \times 4)\]

razor99 (razor99):

u are wrong go to this link

myininaya (myininaya):

I agree with the order of operations which is what razor used 8 + g ÷ h • 4 g=9, h=3 8+9\3 *4 Order of operations tells use to do the dividion/multiplication as it occurs left to right so we have 8+3*4 8+12 Now lastly addition/subtraction (no subtraction here though so really not worth mentioning) 20